# Interpretability-Guided Bi-objective Optimization: Aligning Accuracy and Explainability

### v1

@@ A-A1 Preliminary Definitions
A-A2 Lemma 1: Limitation of Naive Convex Combination
@@
$(\mathbf{g}_{1}\cdot\mathbf{g}_{2})(\|\mathbf{g}_{2}\|^{2}-(\mathbf{g}_{1}\cdot\mathbf{g}_{2}))<\|\mathbf{g}_{2}\|^{2}(\|\mathbf{g}_{1}\|^{2}-(\mathbf{g}_{1}\cdot\mathbf{g}_{2})).$
@@
$(\mathbf{g}_{1}\cdot\mathbf{g}_{2})\|\mathbf{g}_{2}\|^{2}+(\mathbf{g}_{1}\cdot\mathbf{g}_{2})^{2}<\|\mathbf{g}_{1}\|^{2}\|\mathbf{g}_{2}\|^{2}-\|\mathbf{g}_{2}\|^{2}(\mathbf{g}_{1}\cdot\mathbf{g}_{2}).$
@@
A-A3 Theorem 1: Aligned Gradients Case
@@
A-A4 Theorem 2: Conflicting Gradients Case
@@
A-A5 Corollary 1: Descent Property
@@
From Theorems A-A3 and A-A4, we have $\mathcal{P}(\mathbf{g}_{1},\mathbf{g}_{2},\lambda)\cdot\mathbf{g}_{1}>0$. Set $C_{1}(\lambda)=\mathcal{P}(\mathbf{g}_{1},\mathbf{g}_{2},\lambda)\cdot\mathbf{g}_{1}>0$. Similarly for $\mathcal{J}_{2}$ with $C_{2}(\lambda)=\mathcal{P}(\mathbf{g}_{1},\mathbf{g}_{2},\lambda)\cdot\mathbf{g}_{2}>0$. ∎
A-B Proofs for Noise Robustness Properties
A-B1 Theorem 3: Robustness to Gradient Noise
Claim: Let $\tilde{\mathbf{g}}_{1}=\mathbf{g}_{1}+\boldsymbol{\epsilon}_{1}$, $\tilde{\mathbf{g}}_{2}=\mathbf{g}_{2}+\boldsymbol{\epsilon}_{2}$ be noisy gradient estimates with $\mathbb{E}[\boldsymbol{\epsilon}_{i}]=0$ and $Var(\boldsymbol{\epsilon}_{i})=\sigma_{i}^{2}I$. Then the expected update direction $\mathbb{E}[\mathcal{P}(\tilde{\mathbf{g}}_{1},\tilde{\mathbf{g}}_{2},\lambda)]$ is $\mathcal{P}(\mathbf{g}_{1},\mathbf{g}_{2},\lambda)+O(\sigma^{2})$, and the variance is $O(\sigma^{2})$.
@@ A-B1 Theorem 3: Robustness to Gradient Noise
Taking expectation, the linear terms vanish since $\mathbb{E}[\boldsymbol{\epsilon}_{i}]=0$, leaving bias of order $\sigma^{2}$. The variance comes from the linear terms: $Var\left(\frac{\partial\mathcal{P}}{\partial\mathbf{g}_{i}}\boldsymbol{\epsilon}_{i}\right)=O(\sigma_{i}^{2})$. ∎
A-B2 Corollary 2: Effect of Batch Size
@@
For i.i.d. samples, the variance of the gradient estimate scales inversely with batch size: $\sigma^{2}=\sigma_{0}^{2}/B$. Substituting into Theorem A-B1 gives the result. ∎

### v2

@@ V-B1 Preliminary Definitions
V-B2 Lemma 1 ”Limitation of Naive Convex Combination”
@@
$-(\mathbf{g}_{1}\cdot\mathbf{g}_{2})(\|\mathbf{g}_{2}\|^{2}-(\mathbf{g}_{1}\cdot\mathbf{g}_{2}))<\|\mathbf{g}_{2}\|^{2}(\|\mathbf{g}_{1}\|^{2}-(\mathbf{g}_{1}\cdot\mathbf{g}_{2})).$
@@
$-(\mathbf{g}_{1}\cdot\mathbf{g}_{2})\|\mathbf{g}_{2}\|^{2}+(\mathbf{g}_{1}\cdot\mathbf{g}_{2})^{2}<\|\mathbf{g}_{1}\|^{2}\|\mathbf{g}_{2}\|^{2}-\|\mathbf{g}_{2}\|^{2}(\mathbf{g}_{1}\cdot\mathbf{g}_{2}).$
@@
V-B3 Theorem 1 ”Aligned Gradients Case”
@@
V-B4 Theorem 2 ”Conflicting Gradients Case”
@@
V-B5 Corollary 1 ”Descent Property”
@@
From Theorems V-B3 and V-B4, we have $\mathcal{P}(\mathbf{g}_{1},\mathbf{g}_{2},\lambda)\cdot\mathbf{g}_{1}>0$. Set $C_{1}(\lambda)=\mathcal{P}(\mathbf{g}_{1},\mathbf{g}_{2},\lambda)\cdot\mathbf{g}_{1}>0$. Similarly for $\mathcal{J}_{2}$ with $C_{2}(\lambda)=\mathcal{P}(\mathbf{g}_{1},\mathbf{g}_{2},\lambda)\cdot\mathbf{g}_{2}>0$. ∎
V-C Noise Robustness Properties
V-C1 Theorem 3 ”Robustness to Gradient Noise”
Claim: Let $\tilde{\mathbf{g}}_{1}=\mathbf{g}_{1}+\boldsymbol{\epsilon}_{1}$, $\tilde{\mathbf{g}}_{2}=\mathbf{g}_{2}+\boldsymbol{\epsilon}_{2}$ be noisy gradient estimates with $\mathbb{E}[\boldsymbol{\epsilon}_{i}]=0$ and $\mathrm{Var}(\boldsymbol{\epsilon}_{i})=\sigma_{i}^{2}I$. Then the expected update direction $\mathbb{E}[\mathcal{P}(\tilde{\mathbf{g}}_{1},\tilde{\mathbf{g}_{2}},\lambda)]$ is $\mathcal{P}(\mathbf{g}_{1},\mathbf{g}_{2},\lambda)+O(\sigma^{2})$, and the variance is $O(\sigma^{2})$.
@@ V-C1 Theorem 3 ”Robustness to Gradient Noise”
Taking expectation, the linear terms vanish since $\mathbb{E}[\boldsymbol{\epsilon}_{i}]=0$, leaving bias of order $\sigma^{2}$. The variance comes from the linear terms: $\mathrm{Var}\left(\frac{\partial\mathcal{P}}{\partial\mathbf{g}_{i}}\boldsymbol{\epsilon}_{i}\right)=O(\sigma_{i}^{2})$. ∎
V-C2 Corollary 2 ”Effect of Batch Size”
@@
For i.i.d. samples, the variance of the gradient estimate scales inversely with batch size: $\sigma^{2}=\sigma_{0}^{2}/B$. Substituting into Theorem V-C1 gives the result. ∎
Discussion
IGBO’s framework for encoding domain knowledge via DAG constraints opens several interesting research directions. First, the DAG construction methodology naturally accommodates human expert input when gradient information is unavailable. In deployment scenarios where models are treated as black-box computational functions, expert-provided feature importance hierarchies can directly instantiate the DAG, making IGBO applicable even when model internals are inaccessible. Second, the interpretability DAG extracted from a well-performing model provides a blueprint for task-specific model compression. By identifying which feature relationships are essential for a particular task, IGBO can guide the design of smaller, specialized models that retain performance while eliminating unnecessary complexity. This represents a shift from post-hoc explanation toward proactive model design informed by interpretability constraints.
VI Conclusion and Future Work
We have introduced IGBO, a novel framework for training machine learning models to adhere to domain-specific interpretability constraints. By formalizing these constraints via a DAG acting on a differentiable feature importance function $H_{k}(\mathbf{X},\theta)$, and by deriving a geometrically motivated update rule that dynamically decomposes task and interpretability gradients, IGBO provides a principled method to navigate the accuracy-interpretability trade-off. Theoretical analysis proves convergence to Pareto-stationary points via projection mapping $\mathcal{P}$ (Theorems V-B3, V-B4) and establishes robustness to gradient noise (Theorem V-C1). Our CLT-based graph construction provides statistical guarantees on DAG acyclicity and transitivity.
The empirical evaluation demonstrated IGBO’s ability to enforce complex DAG constraints with less than 5% relative accuracy degradation compared to standard training, while achieving over 80% DAG satisfaction rates. The Optimal Path Oracle successfully generated in-distribution integration paths, improving attribution consistency by over 20% compared to linear paths.
Limitations and Future Work:
  • DAG Specification and Human Expertise: The current framework requires the DAG and intervals $\epsilon,\delta]$ to be specified by an expert. In scenarios where gradient access is limited or models are treated as black-box computational functions, human expert annotations can directly provide feature importance rankings. Future work could explore using pairwise comparison data from domain experts to construct the DAG, bypassing gradient computation entirely. This human-in-the-loop approach aligns IGBO with real-world deployment where interpretability constraints often originate from expert knowledge rather than model internals.
  • Task-Specialized Model Compression: Beyond interpretability itself, IGBO enables a pathway toward task-specific model optimization. Given a large model that performs well on a particular task, the structured interpretability DAG extracted under IGBO captures task-relevant feature dependencies. This structure can guide the training of a smaller, task-specialized model that preserves performance while discarding unnecessary capacity. Unlike parameter-efficient fine-tuning approaches that adapt large models in-place, this perspective emphasizes constructing compact models optimized explicitly for a single task, using interpretability as a mechanism for structured knowledge transfer.